Show that the minimum and maximum points of every curve in the family of polynomials 

𝑓(𝑥) = 2𝑥3 + 𝑐𝑥2 + 2𝑥 lie on the curve 𝑦 = 𝑥 − 𝑥3. 


1
Expert's answer
2022-04-27T14:32:09-0400
"f'(x)=6x^2+2cx+2"

Find the critical number(s)


"f'(x)=0=>6x^2+2cx+2=0"

"3x^2+cx+1=0"


"x=\dfrac{-c\pm\sqrt{c^2-12}}{6}"

We consider "x\in \R"


"c^2-12\ge0=>c\le-\sqrt{12}\ or\ c\ge\sqrt{12}"

In this case


"x\not=0, cx^2=-3x^3-x"

Substitute


"y = 2x^3-3x^3-x+2x"

"y = -x^3+x"

Then the minimum and maximum points of every curve in the family of polynomials "f(x)=2x^3+cx^2+2x" lie on the curve "y=x-x^3."

"x\not=0, c\le-\sqrt{12}" or "c\ge\sqrt{12}."